Simplify each of the following to a single logarithm:
1/2log(4x^2)+5log3x-log2

Answer by nerdybill(6958)

1/2log(4x^2)+5log3x-log2
= log(4x^2)^(1/2) + log(3x)^5 - log 2
= log(2x) + log(243x^5) - log 2
= log(2x)(243x^5) - log 2
= log(486x^6) - log 2
= log(486x^6)/2
= log(243x^6)